OpenStudy (anonymous):

5h+2(11•h)=-5 explain please and I'm in algebra I

OpenStudy (saifoo.khan):

5h + 2(11h) = -5 5h + 22h =-5

OpenStudy (saifoo.khan):

27h = -5 h = -5/27

OpenStudy (anonymous):

5h+h=37, please help